a force with magnitude 30n acts on a 1.2 kg soccer ball moving initially in the direction of the force with a speed of 4 m/s. over what distance must the force act to change the ball's speed to 6 m/s?


1
Expert's answer
2021-05-12T08:37:52-0400
"v_f^2-v_i^2=2as=2\frac{F}{m}s\\
6^2-4^2=2\frac{30}{1.2}s\\s=0.4\ m"